# ARM: Refining Multivariate Forecasting with Adaptive Temporal-Contextual Learning

### v1

@@ 4.1 Ablation Studies
Table 2: The results of applying the training strategies in ARM to other LTSF models. We have substituted the encoder-decoder part with various other models for experimentation, testing the effectiveness of transferring AUEL, Random Dropping, and MKLS components of ARM (denoted as A, R, and M, respectively), to other models. Input lengths $L_{I}$ are fixed to 720 for all the experiments. The overall best results for each row of the experiments are highlighted bold. When comparing a model with or without A/R/M modification, the better results are underlined. Note that in part 3, since the sole use of Random Dropping does not impact the performance of univariate models, these results are not provided.
1. ($\pm$ ARM) Vanilla Vanilla+ARM PatchTST PatchTST+ARM DLinear DLinear+ARM Autoformer Autoformer+ARM Informer Informer+ARM
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity 96 0.360 0.434 0.125 0.222 0.137 0.236 0.130 0.225 0.140 0.237 0.129 0.227 0.337 0.423 0.132 0.231 0.922 0.791 0.137 0.238
192 0.377 0.439 0.142 0.239 0.152 0.249 0.145 0.241 0.153 0.249 0.143 0.242 0.310 0.399 0.147 0.241 0.952 0.792 0.149 0.251
336 0.359 0.430 0.154 0.251 0.168 0.265 0.161 0.257 0.169 0.267 0.158 0.262 0.329 0.417 0.156 0.256 0.954 0.796 0.158 0.261
720 0.388 0.436 0.179 0.275 0.207 0.298 0.191 0.293 0.203 0.301 0.187 0.291 0.325 0.419 0.190 0.289 0.941 0.793 0.179 0.276
ETTm1 96 0.917 0.710 0.287 0.340 0.297 0.347 0.285 0.342 0.307 0.351 0.289 0.341 0.544 0.497 0.288 0.346 0.848 0.666 0.293 0.347
192 1.142 0.793 0.328 0.364 0.332 0.370 0.328 0.365 0.339 0.371 0.327 0.363 0.593 0.517 0.329 0.370 0.910 0.702 0.330 0.367
336 1.009 0.739 0.364 0.384 0.367 0.389 0.362 0.387 0.369 0.386 0.365 0.384 0.516 0.482 0.363 0.389 0.977 0.735 0.370 0.391
720 1.211 0.876 0.411 0.412 0.416 0.420 0.414 0.417 0.425 0.421 0.413 0.412 0.554 0.519 0.422 0.421 1.067 0.776 0.421 0.420
2. ($\pm$ A) Vanilla Vanilla+A PatchTST PatchTST+A DLinear DLinear+A Autoformer Autoformer+A Informer Informer+A
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity 96 0.360 0.434 0.130 0.228 0.137 0.236 0.131 0.228 0.140 0.237 0.133 0.232 0.337 0.423 0.131 0.231 0.922 0.791 0.139 0.242
192 0.377 0.439 0.148 0.245 0.152 0.249 0.147 0.242 0.153 0.249 0.150 0.245 0.310 0.399 0.148 0.246 0.952 0.792 0.151 0.252
336 0.359 0.430 0.166 0.264 0.168 0.265 0.163 0.258 0.169 0.267 0.165 0.262 0.329 0.417 0.158 0.259 0.954 0.796 0.159 0.262
720 0.388 0.436 0.187 0.294 0.207 0.298 0.201 0.291 0.204 0.301 0.203 0.294 0.325 0.419 0.197 0.291 0.941 0.793 0.185 0.293
ETTm1 96 0.917 0.710 0.298 0.357 0.297 0.347 0.291 0.343 0.307 0.351 0.300 0.348 0.544 0.497 0.304 0.355 0.848 0.666 0.302 0.354
192 1.142 0.793 0.330 0.367 0.332 0.370 0.332 0.367 0.339 0.371 0.337 0.370 0.593 0.517 0.338 0.374 0.910 0.702 0.336 0.372
336 1.009 0.739 0.369 0.388 0.367 0.389 0.363 0.387 0.369 0.386 0.367 0.383 0.516 0.482 0.370 0.393 0.977 0.735 0.376 0.397
720 1.211 0.876 0.427 0.422 0.416 0.420 0.415 0.417 0.425 0.422 0.423 0.419 0.554 0.519 0.429 0.423 1.067 0.776 0.432 0.425
3. ($\pm$ R) Vanilla Vanilla+R PatchTST PatchTST+R DLinear DLinear+R Autoformer Autoformer+R Informer Informer+R
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity 96 0.360 0.434 0.341 0.416 - - - - - - - - 0.337 0.423 0.207 0.320 0.922 0.791 0.402 0.483
192 0.377 0.439 0.349 0.421 - - - - - - - - 0.310 0.399 0.209 0.319 0.952 0.792 0.360 0.454
336 0.359 0.430 0.348 0.419 - - - - - - - - 0.329 0.417 0.228 0.337 0.954 0.796 0.410 0.488
720 0.388 0.436 0.334 0.407 - - - - - - - - 0.325 0.419 0.316 0.412 0.941 0.793 0.366 0.447
ETTm1 96 0.917 0.710 0.744 0.611 - - - - - - - - 0.544 0.497 0.431 0.458 0.848 0.666 0.458 0.460
192 1.142 0.793 0.973 0.753 - - - - - - - - 0.593 0.517 0.488 0.468 0.910 0.702 0.529 0.521
336 1.009 0.739 0.872 0.674 - - - - - - - - 0.516 0.482 0.481 0.470 0.977 0.735 0.581 0.540
720 1.211 0.876 0.873 0.698 - - - - - - - - 0.554 0.519 0.540 0.512 1.067 0.776 0.697 0.613
4. ($\pm$ M) Vanilla Vanilla+M PatchTST PatchTST+M DLinear DLinear+M Autoformer Autoformer+M Informer Informer+M
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity 96 0.360 0.434 0.378 0.453 0.137 0.236 0.142 0.252 0.140 0.237 0.147 0.256 0.337 0.423 0.221 0.332 0.922 0.791 0.832 0.724
192 0.377 0.439 0.386 0.457 0.152 0.249 0.163 0.269 0.153 0.249 0.165 0.279 0.310 0.399 0.245 0.353 0.952 0.792 0.982 0.775
336 0.359 0.430 0.406 0.468 0.168 0.265 0.173 0.283 0.169 0.267 0.201 0.317 0.329 0.417 0.257 0.362 0.954 0.796 0.730 0.643
720 0.388 0.436 0.365 0.428 0.207 0.298 0.205 0.304 0.203 0.301 0.228 0.338 0.325 0.419 0.272 0.379 0.941 0.793 0.938 0.787
ETTm1 96 0.917 0.710 0.834 0.649 0.297 0.347 0.295 0.344 0.307 0.351 0.305 0.348 0.544 0.497 0.563 0.489 0.848 0.666 0.809 0.623
192 1.142 0.793 0.959 0.752 0.332 0.370 0.339 0.375 0.339 0.371 0.344 0.373 0.593 0.517 0.636 0.513 0.910 0.702 0.903 0.705
336 1.009 0.739 1.005 0.712 0.367 0.389 0.370 0.391 0.369 0.386 0.376 0.394 0.516 0.482 0.567 0.503 0.977 0.735 0.993 0.732
720 1.211 0.876 1.087 0.790 0.416 0.420 0.421 0.419 0.425 0.421 0.435 0.428 0.554 0.519 0.671 0.568 1.067 0.776 1.049 0.775
5. ($\pm$ RM) Vanilla Vanilla+RM PatchTST PatchTST+RM DLinear DLinear+RM Autoformer Autoformer+RM Informer Informer+RM
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity 96 0.360 0.434 0.296 0.384 0.137 0.236 0.131 0.233 0.140 0.237 0.131 0.230 0.337 0.423 0.200 0.315 0.922 0.791 0.272 0.386
192 0.377 0.439 0.294 0.385 0.152 0.249 0.147 0.248 0.153 0.249 0.148 0.245 0.310 0.399 0.199 0.313 0.952 0.792 0.287 0.393
336 0.359 0.430 0.283 0.374 0.168 0.265 0.161 0.261 0.169 0.267 0.164 0.262 0.329 0.417 0.214 0.327 0.954 0.796 0.303 0.377
720 0.388 0.436 0.296 0.397 0.207 0.298 0.197 0.295 0.203 0.301 0.192 0.293 0.325 0.419 0.268 0.372 0.941 0.793 0.331 0.413
ETTm1 96 0.917 0.710 0.713 0.579 0.297 0.347 0.289 0.344 0.307 0.351 0.304 0.346 0.544 0.497 0.272 0.356 0.848 0.666 0.397 0.454
192 1.142 0.793 0.749 0.595 0.332 0.370 0.332 0.369 0.339 0.371 0.338 0.369 0.593 0.517 0.289 0.371 0.910 0.702 0.445 0.467
336 1.009 0.739 0.758 0.610 0.367 0.389 0.364 0.388 0.369 0.386 0.364 0.383 0.516 0.482 0.384 0.407 0.977 0.735 0.541 0.528
720 1.211 0.876 0.815 0.665 0.416 0.420 0.415 0.418 0.425 0.421 0.422 0.420 0.554 0.519 0.443 0.435 1.067 0.776 0.662 0.603
@@
Effects of AUEL   In the second part of Table 2, we showcase the effects of adding the AUEL module to various models. Firstly, AUEL consistently enhances performance, especially for multivariate models: existing multivariate models mixed multiple input series directly without modeling individual univariate effect. In contrast, for univariate models, AUEL’s improvements arise from its adaptive distribution handling. Comparing results from two multivariate models and two univariate models, on the Electricity dataset, where inter-series dependencies are more evident and series follow similar temporal patterns, AUEL-enhanced multivariate models outperform univariate ones. This is attributed to the better recognition of the multivariate relationship after extracting the univariate effect. However, for ETTm1, where series discrepancies are more substantial, making inter-series dependencies learning more challenging, univariate models remain advantageous.
@@
Effects of Random Dropping   In the third of Table 2, we demonstrate the performance gains of multivariate models from the Random Dropping strategy. Since univariate models train each series separately, they are not affected by this strategy, thus not included. Integrating Random Dropping considerably improves the performance of predictors without noticeably increasing computational costs. The improvements are more pronounced in datasets like ETT with weaker inter-series dependencies due to more overfitting reduction. As these relationships strengthen, evident in datasets like Electricity and even more so in Multi, the benefits of Random Dropping tend to diminish.
@@
Effects of MKLS   In the fourth part of Table 2, we show the effects of integrating MKLS into various models. For multivariate input models, MKLS helps learn local patterns across series, particularly evident on datasets with pronounced inter-series relationships like Electricity. For univariate models, adding MKLS is to attempt building multivariate relationship at the token representation level, which can lead to overfitting when series-wise relationships are weak, as shown in the results.
@@
Effects of MKLS with Random Dropping   In the fifth part of Table 2, we highlight the benefits of jointly employing MKLS and Random Dropping. Using both significantly enhances performance compared to their individual applications. Adding Random Dropping reduces potential overfitting during multi-window convolutional layer training and channel-wise attention calculations in MKLS. This synergy allows the MKLS+predictor combination to be trained more effectively for multivariate input. Coupled with AUEL, ARM consistently tackle diverse multivariate forecasting challenges.
@@
5 Conclusion and Future Works
In this study, we introduce ARM, a methodology designed for correctly training multivariate LTSF models. Based on three primary modules: AUEL, Random Dropping, and MKLS, ARM addresses the challenge of handling time series with significant characteristic differences. It achieves this by extracting univariate effects, building reasonable multivariate representations, and discerning effective series combinations for forecasting contribution. Remarkably, with only a minor computational increase, ARM empowers a vanilla Transformer to achieve SOTA performance in LTSF tasks. Each module within ARM can be easily integrated into other LTSF models to enhance their performance, underscoring its practical utility and potential to influence future methodology research in the field.
Furthermore, for future research, potential applications of ARM modules in other time series tasks can be explored. Given that MKLS and Random Dropping decouple relationships between series, it would be worth investigating their viability in time series classification and anomaly detection. Additionally, as AUEL effectively extracts a series’ intrinsic effects, enhancing the extraction of inter-series relationships, its application in tasks like time series clustering can be further explored.
@@ A.2 To What Extent Do We Need Independence in LTSF: A Multivariate Dataset with Significant Dependencies
A.3 Additonal Ablation Studies

### v2

@@ 4.1 Ablation Studies
Table 2: The results of applying A/R/M to more LTSF models. We run all the experiments with fixed $L_{I}=720$. Using percentage comparison, We demonstrated the enhancements in average performance of the original models with the addition of A/R/M. Note that for “+R”, since the sole use of Random Dropping does not impact univariate models, these results are not provided.
Models Vanilla Vanilla+ARM Vanilla+A Vanilla+R Vanilla+M Vanilla+RM
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity (96) 0.360 0.434 0.125 0.222 0.130 0.228 0.341 0.416 0.378 0.453 0.296 0.384
Electricity (192) 0.377 0.439 0.142 0.239 0.148 0.245 0.349 0.421 0.386 0.457 0.294 0.385
Electricity (336) 0.359 0.430 0.154 0.251 0.166 0.264 0.348 0.419 0.406 0.468 0.283 0.374
Electricity (720) 0.388 0.436 0.179 0.275 0.187 0.294 0.344 0.407 0.365 0.428 0.296 0.397
ETTm1 (96) 0.917 0.710 0.287 0.340 0.298 0.357 0.744 0.611 0.834 0.649 0.713 0.579
ETTm1 (192) 1.142 0.793 0.328 0.364 0.330 0.367 0.973 0.753 0.959 0.752 0.749 0.595
ETTm1 (336) 1.009 0.739 0.364 0.384 0.369 0.388 0.872 0.674 1.005 0.712 0.758 0.610
ETTm1 (720) 1.211 0.876 0.411 0.412 0.427 0.422 0.873 0.698 1.087 0.790 0.815 0.665
Average 0.720 0.607 0.249 0.311 0.257 0.321 0.606 0.550 0.678 0.589 0.526 0.499
Average% 100.0% 100.0% 34.6% 51.2% 35.7% 52.9% 84.2% 90.6% 94.2% 97.0% 73.1% 82.2%
Models PatchTST PatchTST+ARM PatchTST+A PatchTST+R PatchTST+M PatchTST+RM
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity (96) 0.137 0.236 0.130 0.225 0.131 0.228 - - 0.142 0.252 0.131 0.233
Electricity (192) 0.152 0.249 0.145 0.241 0.147 0.242 - - 0.163 0.269 0.147 0.248
Electricity (336) 0.168 0.265 0.161 0.257 0.163 0.258 - - 0.173 0.283 0.161 0.261
Electricity (720) 0.207 0.298 0.191 0.293 0.201 0.291 - - 0.205 0.304 0.197 0.295
ETTm1 (96) 0.297 0.347 0.285 0.342 0.291 0.343 - - 0.295 0.344 0.289 0.344
ETTm1 (192) 0.332 0.370 0.328 0.365 0.332 0.367 - - 0.339 0.375 0.332 0.369
ETTm1 (336) 0.367 0.389 0.362 0.387 0.363 0.387 - - 0.370 0.391 0.364 0.388
ETTm1 (720) 0.416 0.420 0.414 0.417 0.415 0.417 - - 0.421 0.419 0.415 0.418
Average 0.260 0.322 0.252 0.316 0.255 0.317 - - 0.264 0.330 0.255 0.320
Average% 100.0% 100.0% 96.9% 98.1% 98.1% 98.4% - - 101.5% 102.5% 98.1% 99.4%
Models DLinear DLinear+ARM DLinear+A DLinear+R DLinear+M DLinear+RM
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity (96) 0.140 0.237 0.129 0.227 0.133 0.232 - - 0.147 0.256 0.131 0.230
Electricity (192) 0.153 0.249 0.143 0.242 0.150 0.245 - - 0.165 0.279 0.148 0.245
Electricity (336) 0.169 0.267 0.158 0.262 0.165 0.262 - - 0.201 0.317 0.164 0.262
Electricity (720) 0.204 0.301 0.187 0.291 0.203 0.294 - - 0.228 0.338 0.192 0.293
ETTm1 (96) 0.307 0.351 0.289 0.341 0.300 0.348 - - 0.305 0.348 0.304 0.346
ETTm1 (192) 0.339 0.371 0.327 0.363 0.337 0.370 - - 0.344 0.373 0.338 0.369
ETTm1 (336) 0.369 0.386 0.365 0.384 0.367 0.383 - - 0.376 0.394 0.364 0.383
ETTm1 (720) 0.425 0.421 0.413 0.412 0.423 0.419 - - 0.435 0.428 0.422 0.420
Average 0.263 0.323 0.251 0.315 0.260 0.319 - - 0.275 0.342 0.258 0.319
Average% 100.0% 100.0% 95.4% 97.5% 98.9% 98.8% - - 104.6% 105.9% 98.1% 98.8%
Models Autoformer Autoformer+ARM Autoformer+A Autoformer+R Autoformer+M Autoformer+RM
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity (96) 0.337 0.423 0.132 0.231 0.131 0.231 0.207 0.320 0.221 0.332 0.200 0.315
Electricity (192) 0.310 0.399 0.147 0.241 0.148 0.246 0.209 0.319 0.245 0.353 0.199 0.313
Electricity (336) 0.329 0.417 0.156 0.256 0.158 0.259 0.228 0.337 0.257 0.362 0.214 0.327
Electricity (720) 0.325 0.419 0.190 0.289 0.197 0.291 0.316 0.412 0.272 0.379 0.268 0.372
ETTm1 (96) 0.544 0.497 0.288 0.346 0.304 0.355 0.431 0.458 0.563 0.489 0.272 0.356
ETTm1 (192) 0.593 0.517 0.329 0.370 0.338 0.374 0.488 0.468 0.636 0.513 0.289 0.371
ETTm1 (336) 0.516 0.482 0.363 0.389 0.370 0.393 0.481 0.470 0.567 0.503 0.384 0.407
ETTm1 (720) 0.554 0.519 0.422 0.421 0.429 0.423 0.540 0.512 0.671 0.568 0.443 0.435
Average 0.439 0.459 0.253 0.318 0.259 0.322 0.363 0.412 0.429 0.437 0.284 0.362
Average% 100.0% 100.0% 57.6% 69.3% 59.0% 70.2% 82.7% 89.8% 97.7% 95.2% 64.7% 78.9%
Models Informer Informer+ARM Informer+A Informer+R Informer+M Informer+RM
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity (96) 0.922 0.791 0.137 0.238 0.139 0.242 0.402 0.483 0.832 0.724 0.272 0.386
Electricity (192) 0.952 0.792 0.149 0.251 0.151 0.252 0.360 0.454 0.982 0.775 0.287 0.393
Electricity (336) 0.954 0.796 0.158 0.261 0.159 0.262 0.410 0.488 0.730 0.643 0.303 0.377
Electricity (720) 0.941 0.793 0.179 0.276 0.185 0.293 0.366 0.447 0.938 0.787 0.331 0.413
ETTm1 (96) 0.848 0.666 0.293 0.347 0.302 0.354 0.458 0.460 0.809 0.623 0.397 0.454
ETTm1 (192) 0.910 0.702 0.330 0.367 0.336 0.372 0.529 0.521 0.903 0.705 0.445 0.467
ETTm1 (336) 0.977 0.735 0.370 0.391 0.376 0.397 0.581 0.540 0.993 0.732 0.541 0.528
ETTm1 (720) 1.067 0.776 0.421 0.420 0.432 0.425 0.697 0.613 1.049 0.775 0.662 0.603
Average 0.946 0.756 0.255 0.319 0.260 0.325 0.475 0.501 0.905 0.721 0.405 0.453
Average% 100.0% 100.0% 27.0% 42.2% 27.5% 43.0% 50.2% 66.3% 95.7% 95.4% 42.8% 59.9%
@@
Effects of AUEL In the third part of Table 2, we showcase the effects of adding the AUEL module to various models. Firstly, AUEL consistently enhances performance, especially for multivariate models: existing multivariate models mixed multiple input series directly without modeling individual univariate effect. In contrast, for univariate models, AUEL’s improvements arise from its adaptive distribution handling. Comparing results from two multivariate models and two univariate models, on the Electricity dataset, where inter-series dependencies are more evident and series follow similar temporal patterns, AUEL-enhanced multivariate models outperform univariate ones. This is attributed to the better recognition of the multivariate relationship after extracting the univariate effect. However, for ETTm1, where series discrepancies are more substantial, making inter-series dependencies learning more challenging, univariate models remain advantageous.
@@
Effects of Random Dropping In the fourth of Table 2, we demonstrate the performance gains of multivariate models from the Random Dropping strategy. Since univariate models train each series separately, they are not affected by this strategy, thus not included. Integrating Random Dropping considerably improves the performance of predictors without noticeably increasing computational costs. The improvements are more pronounced in datasets like ETT with weaker inter-series dependencies due to more overfitting reduction. As these relationships strengthen, evident in datasets like Electricity and even more so in Multi, the improvement of Random Dropping tend to be smaller.
@@
Effects of MKLS In the fifth part of Table 2, we show the effects of integrating MKLS into various models. For multivariate input models, MKLS helps learn local patterns across series, particularly evident on datasets with pronounced inter-series relationships like Electricity. For univariate models, adding MKLS is to attempt building multivariate relationship at the token representation level, which can lead to overfitting when series-wise relationships are weak, as shown in the results.
@@
Effects of MKLS with Random Dropping In the sixth part of Table 2, we highlight the benefits of jointly employing MKLS and Random Dropping. Using both significantly enhances performance compared to their individual applications. Adding Random Dropping reduces potential overfitting during multi-window convolutional layer training and channel-wise attention calculations in MKLS. This synergy allows the MKLS+predictor combination to be trained more effectively for multivariate input. Coupled with AUEL, ARM consistently tackle diverse multivariate forecasting challenges.
@@
5 Conclusion
In this study, we introduce ARM, a methodology designed for effectively training multivariate LTSF models. Based on three primary modules: AUEL, Random Dropping, and MKLS, ARM addresses the challenge of handling time series with significant characteristic differences. It achieves this by extracting univariate effects, building reasonable multivariate representations, and discerning effective series combinations for forecasting contribution. Remarkably, with only a minor computational increase, ARM empowers a vanilla Transformer to achieve SOTA performance in LTSF tasks. Each module within ARM can be easily integrated into other LTSF models to enhance their performance, underscoring its practical utility and potential to influence future methodology research in the field.
@@ A.2 To What Extent Do We Need Independence in LTSF: A Multivariate Dataset with Significant Dependencies
A.3 Additional Explaination of the Three Examples in Figure 1
(a) The necessity of adaptive estimation of output mean for series with different characteristics (see §3.1): We illustrate how different methods estimate the output mean for five example time series. The blue lines denote the original series, with the input and output sections separated by a purple line. Using yellow and red lines, we represent methods that employ the input mean (RevIN) and the last input value (NLinear) respectively as the output mean. These previous methods are suboptimal as they don’t account for varying lookback lengths required for different series. We highlight the ideal lookback areas with green boxes. The Adaptive EMA in our Adaptive Univariate Effect Learning method, represented by a green line, effectively learns the optimal output mean, thereby handling series with diverse temporal dependencies more accurately.
(b) The necessity of building reasonable multivariate temporal representation (see §3.3), which is demonstrated by five series from the Electricity dataset. A black vertical line is used to indicate the current time step $t$. Their associated temporal relationships and local patterns vary for different series. Thus, adopting the values from $X^{(t)}$ as the representation for time step $t$, as done in previous models, wrongly blends the modeling across series with distinct characteristics. We use green, red, and yellow boxes to depict the optimal scope of local patterns that should be considered for different series at time $t$. To accommodate these distinct patterns, we introduce a multi-window local convolutional module to construct reasonable representation for multivariate inputs.
(c) The Inability of Previous Models to Learn Obvious Inter-Series Dependencies: We illustrate the ineffectiveness of existing multivariate LTSF Transformers using a dataset ”Multi”, where the subsequent three series are simple shifts of the first (details in section A.2). A ”copy-paste” operation should suffice for forecasting, yet previous multivariate models, such as Autoformer, fall short to accomplish this task efficiently. In Figure 6, we visually demonstrate ARM’s enhancement in handling these strong inter-series relationships for LTSF models.
A.4 Additonal Ablation Studies